OpenStudy (anonymous):

an: a sphere and a right cylinder have the same radius and volume.find the radius r in terms of the height h of the cylinder

OpenStudy (anonymous):

a sphere and a right cylinder have the same radius and volume.find the radius r in terms of the height h of the cylinder

OpenStudy (anonymous):

help

ganeshie8 (ganeshie8):

you need two formulas : 1) volume of cylinder 2) volume of sphere

OpenStudy (anonymous):

there aren't any at all though

ganeshie8 (ganeshie8):

volume of cylinder = \(\large \pi r^2 h\) volume of sphere = \(\large \frac{4}{3} \pi r^3\)

ganeshie8 (ganeshie8):

Since you're given both volumes are equal, set them equal and solve \(\large r\)

ganeshie8 (ganeshie8):

set them equal : \(\large \pi r^2 h = \frac{4}{3} \pi r^3\)

ganeshie8 (ganeshie8):

you can cancel pi and r^2 both sides : \(\large h = \frac{4}{3} r\)

OpenStudy (anonymous):

okay

ganeshie8 (ganeshie8):

can you solve \(\large r\) now ?

OpenStudy (anonymous):

should i choose any number

ganeshie8 (ganeshie8):

nope, they just want the \(r\) in terms of \(h\)

ganeshie8 (ganeshie8):

\(\large h = \frac{4}{3} r\) multiply 4/3 both sides \(\large \frac{3}{4} h = r\)

ganeshie8 (ganeshie8):

which is same as \(\large r=\frac{3}{4} h\)

OpenStudy (anonymous):

so i chose 5 first was 25

OpenStudy (anonymous):

so 25 and 17.5?

ganeshie8 (ganeshie8):

from where u getting those numbers :/

ganeshie8 (ganeshie8):

\(\large r = \frac{3}{4} h \) and its the end. they just want u find out \(r\) in terms of \(h\)

ganeshie8 (ganeshie8):

we're done^